OpenStudy (love_to_love_you):

Write an equation for the translation of y = 2/x that has the given asymptotes. x = 4 and y = -8 Show all work

OpenStudy (love_to_love_you):

@zepdrix

jimthompson5910 (jim_thompson5910):

what is the vertical asymptote for y = 2/x ?

OpenStudy (love_to_love_you):

oh nvm I don't even have to do this problem. Could you help me with something else?

OpenStudy (love_to_love_you):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (love_to_love_you):

Okay so what I have to do is explain why for each problem. I have the answers already but I do not know the explanation.

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (love_to_love_you):

a. If sin theta = sqrt 2/2, which could not be the value of theta? 225 degrees

jimthompson5910 (jim_thompson5910):

where is 225 degrees? which quadrant?

OpenStudy (love_to_love_you):

Um is it quadrant 3?

jimthompson5910 (jim_thompson5910):

yes, and sine is negative in Q3 and Q4

jimthompson5910 (jim_thompson5910):

so it's impossible for theta to be 225 degrees

OpenStudy (love_to_love_you):

ok

OpenStudy (love_to_love_you):

b. For which value of theta is tan theta equal to sin theta? 2pi

jimthompson5910 (jim_thompson5910):

I'm guessing you had a list of choices?

OpenStudy (love_to_love_you):

It's a problem. I have 2 problems. They have a and b parts to them

OpenStudy (love_to_love_you):

So what we have to do is explain why this is the correct answer/ show work.

jimthompson5910 (jim_thompson5910):

well tan(theta) is the same as sin(theta)/cos(theta) it's one of the many identities

jimthompson5910 (jim_thompson5910):

\[\Large \tan(\theta) = \sin(\theta)\] \[\Large \frac{\sin(\theta)}{\cos(\theta)} = \sin(\theta)\] \[\Large \frac{\sin(\theta)}{\cos(\theta)}*{\color{red}{\frac{1}{\sin(\theta)}}} = \sin(\theta)*{\color{red}{\frac{1}{\sin(\theta)}}}\] \[\Large \frac{\cancel{\sin(\theta)}}{\cos(\theta)}*{\color{black}{\frac{1}{\cancel{\sin(\theta)}}}} = \cancel{\sin(\theta)}*{\color{black}{\frac{1}{\cancel{\sin(\theta)}}}}\] \[\Large \frac{1}{\cos(\theta)} = 1\] in step 3, I'm multiplying both sides by 1/sin(theta) since both sides have a sine that cancels making sense so far?

OpenStudy (love_to_love_you):

Yeah

OpenStudy (love_to_love_you):

so what happens next?

jimthompson5910 (jim_thompson5910):

multiply both sides by cos(theta) to get cos(theta) = 1 then you'll use arccosine to isolate theta

OpenStudy (love_to_love_you):

Idk how to use arccosine very well

jimthompson5910 (jim_thompson5910):

what kind of calculator do you have?

OpenStudy (love_to_love_you):

just an online graphing calc

jimthompson5910 (jim_thompson5910):

what's the link to it? so I can have a look

OpenStudy (love_to_love_you):

it's a download from my school

jimthompson5910 (jim_thompson5910):

I gotcha

jimthompson5910 (jim_thompson5910):

ok I'm going to use this calculator here http://web2.0calc.com/

OpenStudy (love_to_love_you):

ok

jimthompson5910 (jim_thompson5910):

click the "rad" button (next to "deg") to convert over to radian mode

jimthompson5910 (jim_thompson5910):

then type in "arccos(1)" without quotes http://web2.0calc.com/#arccos(1) what do you get?

OpenStudy (love_to_love_you):

cos^-1(1) = 0

OpenStudy (love_to_love_you):

and then there's a tiny 2 pi under it

OpenStudy (love_to_love_you):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct, so 0 radians is one answer notice how 2pi radians and 0 radians are coterminal angles

jimthompson5910 (jim_thompson5910):

so since theta = 0 is one answer, theta = 2pi is also an answer there are infinitely many other answers. They only want theta = 2pi for some reason

OpenStudy (love_to_love_you):

So that makes 2 pi the answer as well?

jimthompson5910 (jim_thompson5910):

yeah there are infinitely many answers, but the computer or teacher is only accepting 2pi

OpenStudy (love_to_love_you):

ok

OpenStudy (love_to_love_you):

I'm gonna try to figure out a for the next problem on my own but can you help me with b?

OpenStudy (love_to_love_you):

A man stands on his balcony, 140 feet above the ground. He looks at the ground, with his sight line forming an angle of 75 degrees with the building, and sees a bus stop. The function d = 140 sec theta models the distance from the man to any object given his angle of sight theta. How far is the bus stop from the man? Round your answer. 541 ft.

jimthompson5910 (jim_thompson5910):

140*sec(theta) = 140*(1/cos(theta)) plug in theta = 75 make sure you are in degree mode use this calculator if needed http://web2.0calc.com/

OpenStudy (love_to_love_you):

541

jimthompson5910 (jim_thompson5910):

yeah I'm getting 540.9184627208766586 which rounds to 541 (assuming you round to the nearest foot)

OpenStudy (love_to_love_you):

yeah i rounded

OpenStudy (love_to_love_you):

Thanks so much

jimthompson5910 (jim_thompson5910):

you're welcome